OpenStudy (anonymous):

permutations : teacher must arrange 5 tunes for band. she has 20 tunes to choose from. how many arrangements are possible?

OpenStudy (anonymous):

do you know how to compute this using the counting principle?

OpenStudy (anonymous):

you have 20 choices for the first one then 19 for the second 18 for the third 17 for the fourth 16 for the fifth

OpenStudy (anonymous):

can we not do the counting principle?? there's a formula.. nPr

OpenStudy (anonymous):

counting principle says the number of ways to do these together is the product i.e. multiply them

OpenStudy (anonymous):

the formula \(_nP_r\) is just an encapsulation of the counting principle for this type of problem it is based on the counting principle

OpenStudy (anonymous):

don't get married to the formula, just think \[20\times 19\times 18\times 17\times 16\]

OpenStudy (anonymous):

that's not the answer though.. 20

OpenStudy (anonymous):

20!

OpenStudy (anonymous):

the answer is much smaller

OpenStudy (anonymous):

1860480 is answer